OpenStudy (abhisar):

@Abmon98 @aaronq @sidsiddhartha @ParthKohli

OpenStudy (abhisar):

If V in the equation \(\sf \lambda_{eq}=\kappa \times V\), is the volume in CC containing 1 eq. of the electrolyte; V for N/10 solution will be ?

OpenStudy (abhisar):

If the volume of solution containing one gram equivalent of the electrolyte is V \(cm^3\), then Equivalent Conductivity \(\sf (\lambda_eq)=Specific~conductivity~(\kappa)\times V\)

OpenStudy (abhisar):

In terms of concentration it means, if the solution has a concentration of c gram equivalent per liter i.e c gram equivalents are present in 1000 cm3 of the solution, then the volume of the solution containing one gram equivalent will be \(\sf \large \frac{1000}{c}\\ or ~V=\frac{\Large 1000}{\Large c}=\frac{1000}{\LARGE Normality}\) In the question, Normality given= N/10 = 0.1N \(\sf \huge V=\frac{1000}{0.1}=10,000CC\)